# On the Closed-Form of Flow Matching: Generalization Does Not Arise from Target Stochasticity

### v1

@@ Appendix B Additional details and comments on empirical flow matching
First, recalls on the optimal velocity (Equation 6) and the empirical flow matching loss (Equations 7 and 8) are provided in Section B.1. The unbiasedness of the estimator is presented in Section B.2, and its proof is in Section B.3.
@@ B.1 Recalls
$\hat{u}^{\star}(x,t)$ $=\sum_{l=1}^{n}\frac{x^{(l)}-x}{1-t}\cdot\left[\mathrm{softmax}\left(\left(-\frac{\lVert{x-tx^{(k)}}\rVert^{2}}{2(1-t)^{2}}\right)_{k=1,\dots,n}\right)\right]_{l}\enspace.$ (6)
@@
$\mathcal{L}_{\mathrm{EFM}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0}\\ x_{1}\sim\hat{p}_{\mathrm{data}}\\ x_{t}=(1-t)x_{0}+tx_{1}\\ b^{(1)}:=x_{1}\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\enspace,$ (7)
@@
$\hat{u}_{M}^{\star}(x_{t},t)$ $=\sum_{j=1}^{M}\frac{b^{(j)}-x_{t}}{1-t}\cdot\left[\mathrm{softmax}\left(\left(-\frac{\lVert{x_{t}-tb^{(k)}}\rVert^{2}}{2(1-t)^{2}}\right)_{k=1,\dots,M}\right)\right]_{j}\enspace.$ (8)
Crucially, in Equation 7 the sample $b^{(1)}$ depends on $x_{t}$ and is reused in the estimate $\hat{u}_{M}^{\star}$. This important detail yields an unbiased estimator of $\hat{u}^{\star}$.
@@ B.2 Theoretical properties of the proposed estimator
First, we discuss below the relation between Proposition 2 and the sampling literature.
Links with importance sampling. The estimator $\hat{u}^{\star}$ in Equation 6 can be seen as a form of importance sampling (see Robert et al. 1999, Chap. 3 for an in-depth reference). In a nutshell, importance sampling is a way to estimate an expectation when one cannot easily sample from the random variable it depends on. More precisely, in the ideal case $z\sim p_{\mathrm{data}}$ (as opposed to $z\sim\hat{p}_{\mathrm{data}}$), the velocity field formula is the following
@@
$=\int_{z}u^{\mathrm{cond}}(x_{t},z,t)p(z|x_{t},t)\mathrm{d}z\enspace.$ (28)
@@
$\phantom{\approx}:=\hat{u}^{\star}(x_{t},t)\enspace.$ (32)
B.3 Proof of Proposition 2
We first recall Appendix B, which we prove in this section. See 2
@@ B.3 Proof of Proposition 2
$\;\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{1}\sim\hat{p}_{\mathrm{data}}\\ x_{t}=(1-t)x_{0}+tx_{1}\\ b^{(1)}:=x_{1}\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\enspace,$ (33)
=\;= $\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{t}\sim p_{t}\\ \end{subarray}}\mathbb{E}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\\ b^{(2)},\dots,b^{(M)}|x_{t},t\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\enspace,$ (34)
=\;= $\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{t}\sim p_{t}\\ \end{subarray}}\mathbb{E}_{\begin{subarray}{c}b^{(1)}:=\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\enspace\text{because $b^{(2)},\dots,b^{(M)}\perp\!\!\!\perp x_{t},t$}\enspace,$ (35)
@@
$u_{\theta}(x_{t},t)=\mathbb{E}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left[\hat{u}_{M}^{\star}(x_{t},t)\right]\enspace.$ (36)
@@
$\hat{u}^{\star}(x_{t},t)=\hat{u}^{\star}$ $=\sum_{l=1}^{n}w^{(l)}u^{(l)}\enspace,\text{ with }$ (37)
$w^{(l)}$ $=\hat{p}_{\mathrm{data}}(z=x^{(l)}|t,x_{t})\enspace,\;\sum_{l=1}^{n}w^{(l)}=1$ (38)
@@
We express a mini-batch as an $M$-valued vector of indices, $\bm{i}\in\llbracket 1,n\rrbracket^{M}$. The mini-batch estimate from Equation 7, considering the definition of the $\mathrm{softmax}$, can be expressed as a mini-batch weighted-mean:
@@
The main result of the following is that, in expectation over the biased-mini-batches, where the first point is drawn according to $w$ and the $M-1$ other points are drawn uniformly, the mini-batch weighted-mean is an unbiased estimate of the $w$-weighted-mean $\hat{u}^{\star}$.
@@
The expression in Equation 46 is invariant with respect to order of the indices $\bm{i}_{1},\dots,\bm{i}_{M}$: the indices in expectation in Equation 46 can be exchanged, and one thus has
${\color[rgb]{0.0,0.4,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.4,0.13}\forall k\in\llbracket 1,M\rrbracket},\;\mathbb{E}\left[\hat{u}_{M}^{\star}({\bf{i}})\right]=n\;\mathbb{E}_{\bm{i}_{1},...,\bm{i}_{M}\sim\operatorname{Unif}}\left[w^{({{\color[rgb]{0.0,0.4,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.4,0.13}\bm{i}_{k}}})}\hat{u}_{M}^{\star}(\bf{i})\right]\enspace.$ (47)
Averaging Equation 47 over the indices $k\in\llbracket 1,M\rrbracket$ yields the desired result
@@
$\phantom{\leq}=\mathbb{E}_{l\sim\mathrm{Unif}}\left[w^{(l)}(u^{(l)})^{2}\right]\enspace.$ (69)
@@
$\mathbb{E}_{x^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x_{t},t)^{2}\right]-(\hat{u}^{\star})^{2}\leq\mathbb{E}_{l\sim\mathrm{Unif}}\left[w^{(l)}(u^{(l)})^{2}\right]-(\hat{u}^{\star})^{2}\enspace,$ (70)
@@
$\mathrm{Var}_{x^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x_{t},t)\right]\leq\mathrm{Var}_{x^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)}\left[u^{\mathrm{cond}}(x_{t},x^{(1)},t)\right]\enspace.$ (71)
Appendix C Experiments details
For all the experiment we used all the same learning hyperparameters, the default ones form Tong et al. [2024]. The hyperparameter values are summarized in Table 1. The details specific to each figure are described in Sections C.1, C.2, C.3 and C.4
##\## Channels Batch Size Learning Rate EMA Decay Gradient Clipping
@@ Appendix C Experiments details
C.1 Figures 1(b) and 2
@@
For Figure 1(b) no deep learning is involved: the datasets 2-moons and CIFAR-10 are loaded. Then, $256$ points from $p_{0}\times\hat{p}_{\mathrm{data}}$ are drawn, and one computes the mean of the cosine similarities between $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$, for each value of $t\in\{0,1/100,2/100,\dots,99/100\}$.
@@
No deep learning either is involved in Figure 2: the Imagenette dataset is loaded and spatially subsampled to resolution $\mathrm{dim}=8$, $\mathrm{dim}=16$, …, $\mathrm{dim}=256$, i.e., with $d="\cdot 8^{2}$, $d=3\cdot 16^{2}$, …, $d=3\cdot 256^{2}$. Then, as for Figure 1(b), batches of $256$ points from $p_{0}$ and $p_{\mathrm{data}}$ are drawn, and one computes the percentage of cosine similarities between $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$, that are larger than $0.9$, for multiple time values $t$.
C.2 Figure 3
@@
In Figure 3, networks are trained with a vanilla conditional flow matching, with the standard 34 million parameters U-Net for diffusion by Nichol and Dhariwal [2021], with default settings from the torchfm codebase  [^7] https://github.com/atong01/conditional-flow-matching [Tong et al., 2024]. Training uses the CFM loss. For this specific experiment, we removed the usual random flip transform, for $\hat{u}^{\star}$ to be simpler and easier to estimate by $u_{\theta}$. For each “data” subsampling of the dataset, we trained the model for $5\cdot 10^{4}$ iterations, with a batch size of $128$, i.e., we trained the models for $128$ epochs.
C.3 Figure 4

### v2

@@ Appendix B Additional details and comments on empirical flow matching
First, recalls on the optimal velocity (Equation˜6) and the empirical flow matching loss (Equations˜7 and 8) are provided in Section˜B.1. The unbiasedness of the estimator is presented in Section˜B.2, and its proof is in Section˜B.3.
@@ B.1 Recalls
$\hat{u}^{\star}(x,t)$ $=\sum_{l=1}^{n}\frac{x^{(l)}-x}{1-t}\cdot\left[\mathrm{softmax}\left(\left(-\frac{\lVert{x-tx^{(k)}}\rVert^{2}}{2(1-t)^{2}}\right)_{k=1,\dots,n}\right)\right]_{l}\kern 5.0pt.$ (6)
@@
$\mathcal{L}_{\mathrm{EFM}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0}\\ x_{1}\sim\hat{p}_{\mathrm{data}}\\ x_{t}=(1-t)x_{0}+tx_{1}\\ b^{(1)}:=x_{1}\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\kern 5.0pt,$ (7)
@@
$\hat{u}_{M}^{\star}(x_{t},t)$ $=\sum_{j=1}^{M}\frac{b^{(j)}-x_{t}}{1-t}\cdot\left[\mathrm{softmax}\left(\left(-\frac{\lVert{x_{t}-tb^{(k)}}\rVert^{2}}{2(1-t)^{2}}\right)_{k=1,\dots,M}\right)\right]_{j}\kern 5.0pt.$ (8)
Crucially, in Equation˜7 the sample $b^{(1)}$ depends on $x_{t}$ and is reused in the estimate $\hat{u}_{M}^{\star}$. This important detail yields an unbiased estimator of $\hat{u}^{\star}$.
@@ B.2 Theoretical properties of the proposed estimator
First, we discuss below the relation between Proposition˜2 and the sampling literature.
Links with importance sampling. The estimator $\hat{u}^{\star}$ in Equation˜6 can be seen as a form of importance sampling (see robert1999montecarlobook for an in-depth reference). In a nutshell, importance sampling is a way to estimate an expectation when one cannot easily sample from the random variable it depends on. More precisely, in the ideal case $z\sim p_{\mathrm{data}}$ (as opposed to $z\sim\hat{p}_{\mathrm{data}}$), the velocity field formula is the following
@@
$=\int_{z}u^{\mathrm{cond}}(x_{t},z,t)p(z|x_{t},t)\mathrm{d}z\kern 5.0pt.$ (31)
@@
$\phantom{\approx}:=\hat{u}^{\star}(x_{t},t)\kern 5.0pt.$ (35)
B.3 Proof of Proposition˜2
We first recall Appendix˜B, which we prove in this section. See 2
@@ B.3 Proof of Proposition˜2
$\;\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{1}\sim\hat{p}_{\mathrm{data}}\\ x_{t}=(1-t)x_{0}+tx_{1}\\ b^{(1)}:=x_{1}\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\kern 5.0pt,$ (36)
$=\;$ $\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{t}\sim p_{t}\\ \end{subarray}}\mathbb{E}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\\ b^{(2)},\dots,b^{(M)}|x_{t},t\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\kern 5.0pt,$ (37)
$=\;$ $\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{t}\sim p_{t}\\ \end{subarray}}\mathbb{E}_{\begin{subarray}{c}b^{(1)}:=\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\kern 5.0pt\text{because $b^{(2)},\dots,b^{(M)}\perp\!\!\!\perp x_{t},t$}\kern 5.0pt,$ (38)
@@
$u_{\theta}(x_{t},t)=\mathbb{E}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left[\hat{u}_{M}^{\star}(x_{t},t)\right]\kern 5.0pt.$ (39)
@@
$\hat{u}^{\star}(x_{t},t)=\hat{u}^{\star}$ $=\sum_{l=1}^{n}w^{(l)}u^{(l)}\kern 5.0pt,\text{ with }$ (40)
$w^{(l)}$ $=\hat{p}_{\mathrm{data}}(z=x^{(l)}|t,x_{t})\kern 5.0pt,\;\sum_{l=1}^{n}w^{(l)}=1$ (41)
@@
We express a mini-batch as an $M$-valued vector of indices, $\bm{i}\in\llbracket 1,n\rrbracket^{M}$. The mini-batch estimate from Equation˜7, considering the definition of the $\mathrm{softmax}$, can be expressed as a mini-batch weighted-mean:
@@
The main result of the following is that, in expectation over the biased-mini-batches, where the first point is drawn according to $w$ and the M1M-1 other points are drawn uniformly, the mini-batch weighted-mean is an unbiased estimate of the $w$-weighted-mean $\hat{u}^{\star}$.
@@
The expression in Equation˜49 is invariant with respect to order of the indices $\bm{i}_{1},\dots,\bm{i}_{M}$: the indices in expectation in Equation˜49 can be exchanged, and one thus has
${\color[rgb]{0.0,0.4,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.4,0.13}\forall k\in\llbracket 1,M\rrbracket},\;\mathbb{E}\left[\hat{u}_{M}^{\star}({\bf{i}})\right]=n\;\mathbb{E}_{\bm{i}_{1},...,\bm{i}_{M}\sim\operatorname{Unif}}\left[w^{({{\color[rgb]{0.0,0.4,0.13}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.4,0.13}\bm{i}_{k}}})}\hat{u}_{M}^{\star}(\bf{i})\right]\kern 5.0pt.$ (50)
Averaging Equation˜50 over the indices $k\in\llbracket 1,M\rrbracket$ yields the desired result
@@
$\phantom{\leq}=\mathbb{E}_{l\sim\mathrm{Unif}}\left[w^{(l)}(u^{(l)})^{2}\right]\kern 5.0pt.$ (72)
@@
$\mathbb{E}_{x^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x_{t},t)^{2}\right]-(\hat{u}^{\star})^{2}\leq\mathbb{E}_{l\sim\mathrm{Unif}}\left[w^{(l)}(u^{(l)})^{2}\right]-(\hat{u}^{\star})^{2}\kern 5.0pt,$ (73)
@@
$\mathrm{Var}_{x^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x_{t},t)\right]\leq\mathrm{Var}_{x^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x_{t},t)}\left[u^{\mathrm{cond}}(x_{t},x^{(1)},t)\right]\kern 5.0pt.$ (74)
Appendix C Additional experiments
We present below the results for the MNIST dataset. The conclusions atre the same as for the CIFAR-10 and CelebA $64\times 64$: regressing against a more deterministic velocity field does not hurt generalization. On the contrary, generalization (i.e., lower test FID) appears earlier during training.
For this experiment, we used the Unet with attention and timestep embedding from torchcfm library, with the Adam optimizer and all the default parameters. We used a pretrained classifier with $99\%$ accuracy on MNIST ($90\%$ on FMNIST) as a lower-dimensional embedding of size $128$ to compute the FID between the test set and the generated set.
Method Ep. 1 Ep. 2 Ep. 3 Ep. 4 Ep. 5 Ep. 10 Ep. 15 Ep. 20 Ep. 25
CFM (EFM, M=1) 378.00 181.25 67.88 29.44 15.30 4.20 3.08 2.51 2.28
EFM, M=128 370.64 168.58 60.52 25.52 13.44 3.79 2.70 2.35 2.10
EFM, M=256 370.94 169.71 61.88 25.73 13.48 3.73 2.76 2.33 2.08
EFM, M=1024 369.72 168.43 60.28 24.24 12.26 3.30 2.67 2.17 1.84
Table 1: FID FMNIST. FID scores across training epochs for conditional flow matching and empirical flow matching for multiple values of the number of samples $M$ used to estimate the closed-form $\hat{u}^{\star}$.
Method FID Ep. 5 FID Ep. 10 FID Ep. 50 FID Ep. 100 FID Ep. 200
CFM (EFM, M=1) 253.56 48.67 25.36 21.35 19.67
EFM, M=128 206.27 44.08 23.39 19.63 17.72
EFM, M=256 202.62 45.06 22.16 20.08 17.74
EFM, M=512 194.66 44.19 22.10 18.93 16.85
Table 2: FID FMNIST. FID scores across training epochs for conditional flow matching and empirical flow matching for multiple values of the number of samples $M$ used to estimate the closed-form $\hat{u}^{\star}$.
Appendix D Experiments details
For all the experiment we used all the same learning hyperparameters, the default ones form tong2023improving. The hyperparameter values are summarized in Table˜3. The details specific to each figure are described in Sections˜D.2, D.3, D.4 and D.5
#\# Channels Batch Size Learning Rate EMA Decay Gradient Clipping
@@ Appendix D Experiments details
D.1 Compute time
Given that regressing against an estimate of the closed-form, EFM, seems to improve on CFM, one may wonder what is the additional cost induced by EFN. To alleviate the non-linearity of GPU computing (parallelism may cause some discontinuities in terms of costs), we ran an exhaustive set of timing experiments, varying the batch size and the EFM sample size. To summarize the measurements (numbers are given for an NVIDIA L4 GPU, on CIFAR-10), denoting $b$ the batch size and $e$ the EFM sample size, the cost follows $b\times(4.3ms+e\times 0.9\mu s)$. It can be also be seen as adding $\sim 2\%$ for every 100 EFM samples. Or, for instance with a batch size of 256, $1.1$ second will be due to the $256$-sample forward/backward, while the additional cost for EFM-$1000$ will be 230230ms (around $17\%$ of the cost) and for EFM-$128$ under 3030ms (under $3\%$).
D.2 Figures˜1(a) and 1(c)
@@
For Figure˜1(a) no deep learning is involved: the datasets 2-moons and CIFAR-10 are loaded. Then, $256$ points from $p_{0}\times\hat{p}_{\mathrm{data}}$ are drawn, and one computes the mean of the cosine similarities between $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$, for each value of $t\in\{0,1/100,2/100,\dots,99/100\}$.
@@
No deep learning either is involved in Figure˜1(c): the Imagenette dataset is loaded and spatially subsampled to resolution $\mathrm{dim}=8$, $\mathrm{dim}=16$, …, $\mathrm{dim}=256$, i.e., with $d="\cdot 8^{2}$, $d=3\cdot 16^{2}$, …, $d=3\cdot 256^{2}$. Then, as for Figure˜1(a), batches of $256$ points from $p_{0}$ and $p_{\mathrm{data}}$ are drawn, and one computes the percentage of cosine similarities between $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$, that are larger than $0.9$, for multiple time values $t$.
D.3 Figure˜2
@@
In Figure˜2, networks are trained with a vanilla conditional flow matching, with the standard 34 million parameters U-Net for diffusion by nichol2021improved, with default settings from the torchfm codebase  [^3] https://github.com/atong01/conditional-flow-matching [tong2023improving]. Training uses the CFM loss. For this specific experiment, we removed the usual random flip transform, for $\hat{u}^{\star}$ to be simpler and easier to estimate by $u_{\theta}$. For each “data” subsampling of the dataset, we trained the model for $5\cdot 10^{4}$ iterations, with a batch size of $128$, i.e., we trained the models for $128$ epochs.
D.4 Figure˜3